OpenStudy (anonymous):

Express the function y=8(x-8)^5 as a composition y=f(g(x)) of two simpler functions y=f(u) and u=g(x)

OpenStudy (mathteacher1729):

We're given \[\huge y = 8(x-8)^2\] and you can look at it as "eight times (some function) squared" or... \[\huge y = 8\Box^2 \] What is \(\huge \Box \)?

OpenStudy (anonymous):

x-8 ??

OpenStudy (mathteacher1729):

@yoadriiii Yes. So that means \[\huge f(u) = 8u^2\] and \[\huge g(x) = (x-8)\]

OpenStudy (anonymous):

okay and would it be ^2 or ^5 because I posted the power of 5 not 2 lol just making sure ??

OpenStudy (mathteacher1729):

Oh sorry, yeah it should be 5. my eyes. they are going. :-p

OpenStudy (anonymous):

lol it's okay thank you so much forthe help. I actually have another one... if the problem now shows y=3x\[^{6}\]-2

OpenStudy (anonymous):

sorry I inputed that wrong

OpenStudy (anonymous):

The problem is y=3x^6-2 and I again have to find g(x) and f(u)

OpenStudy (anonymous):

I'm confused again because now it doesn't show the parentheses?